if P=sintsin(t+pi/3) then find average value ofP over the interval t=0 to t=2pi.

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

sintsin(t+pi/3)= (1/2)(cos(pi/3)-cos(2t+pi/3)), period of cos(2t+pi/3) is pi, therefore average value of cos(2t+pi/3)=zero,

thus average value of p=(1/2)cos(pi/3)=1/4